OpenStudy (anonymous):

use the equation 2(2-y)-3x. what are the values of y when x= -4 , 0 , 2/3 , and 3

OpenStudy (texaschic101):

x = -4 2(2 - y) - 3x 4 - 2y - 3(-4) 4 - 2y + 12 -2y + 8 x = 0 2(2 - y) - 3(0) 4 - 2y - 0 -2y + 4 x = 2/3 2(2 - y) - 3(2/3) 4 - 2y - 6/3 -2y - 6/3 + 4 -2y -6/3 + 12/3 -2y + 6/3 -2y + 2 x = 3 2(2 - y) - 3(3) 4 - 2y - 9 -2y - 9 + 4 -2y - 5